OpenStudy (anonymous):

The ordered pairs <–8, 2> and <5, –6> represent two vectors. What is the magnitude of the resultant? (1 point)1 2 5 15

jimthompson5910 (jim_thompson5910):

Resultant vector: <–8, 2> + <5, –6> = <-8+5, 2+(-6)> = <-3, -4> Magnitude of resultant vector = sqrt((-3)^2+(-4)^2) = sqrt(9+16)=sqrt(25) = 5

jimthompson5910 (jim_thompson5910):

So the answer is choice B) 5